# Effect of freestream turbulence on the coherent dynamics of a wind turbine wake

### v1

@@ 6 Energy exchanges
The kinetic energy budget terms are first discussed for the four cases analysed previously. The energy budget terms are integrated over the entire domain and shown in figure 16. Since the modes related to the tip vortices show weaker energy exchanges compared to the low frequency modes (particularly for the FST cases), a separate colour bar is used for them. As a general observation, as freestream turbulence level is increased, the number of energetic low frequency modes increased. Further, the number of tip vortex related modes decreased as discussed previously. Biswas and Buxton [2024b] showed that in the presence of negligible freestream turbulence, the wake meandering mode acted like a primary mode, extracting energy from the mean flow. In the presence of FST, the mean flow production ($\tilde{P}_{l}$) term is still found to be the dominant energy source for wake meandering. The $\tilde{P}_{l}$ terms for the dominant wake meandering modes in the presence of FST were always found to be stronger than that for the no grid case which is in line with that discussed in section 4.2. Furthermore, the dominant wake meandering modes for case 2a drew the highest amount of energy from the mean flow. The $\tilde{P}_{l}$ term then became relatively weaker as $T_{i}$ was increased further, likely due to the wake meandering modes becoming less coherent (as observed in figure 13) and less efficient at extracting energy from the mean flow. This was accompanied by an enhancement in the relative contributions from the triadic interaction and diffusion term for the highest $T_{i}$ case.
The intriguing question is why is the $\tilde{P}_{l}$ term at all stronger for the wake meandering modes in the presence of FST? To answer this we need to look at the composition of the $\tilde{P}_{l}$ term. As discussed in Biswas and Buxton [2024b], utilizing the fact that the transverse gradient of streamwise mean velocity ($\frac{\partial\bar{u}}{\partial y}$) is much stronger than the other mean velocity gradients and assuming that the velocity components of the different modes are uncorrelated ($\overline{\tilde{u}^{l}\tilde{v}^{n}}\approx 0$), we can approximate the mean flow production term as $\tilde{P}_{l}\approx\overline{\tilde{u}^{l}\tilde{v}^{l}}\frac{\partial\overline{u}}{\partial y}$. This was verified to be true for all the modes and is not shown here for brevity, a more detailed discussion can be found in Biswas and Buxton [2024b]. In figure 17, the profiles of $\overline{\tilde{u}^{l}\tilde{v}^{l}}$ and $\frac{\partial\overline{u}}{\partial y}$ are shown for the wake meandering mode extracting the highest amount of energy from the mean flow for the no grid case (figure 17(a,b)) and case 2a (figure 17(c,d)), which is representative of all FST cases. These are shown at two different streamwise locations $x/D=1$ and $x/D=3$. The Strouhal numbers of the modes are $St_{D}=0.22$ (no grid case) and $St_{D}=0.37$ (case 2a). Note that, the no grid case shows the strongest gradients in the tip shear layer. However the wake meandering mode remains sheltered well within the shear layer (see figure 13(a)), inhibiting it from utilizing the stronger gradients available at the shear layer. For case 2a, the wake meandering mode extends beyond the shear layer (see figure 13(d)). As a result, although the mean velocity gradients are slightly weaker (due to enhanced turbulent mixing), the mode does a better job of exploiting this shear thereby yielding a higher rate of energy transfer from the mean flow. A similar trend was recorded for the other wake meandering modes with FST showing a higher $\tilde{P}_{l}$ term (not shown here for brevity).
Figure 18: Inter frequency triadic energy transfers for (a) no grid case, (b) case 2a and (c) case 3.
@@
The tip vortex related modes in the presence of FST showed energy exchanges similar to the no grid case, however, the magnitudes of the terms decreased as $T_{i}$ was increased. $f_{r}$ for the no grid case was primarily energized by the mean flow production ($\tilde{P}_{l}$) term. For the FST cases, the relative contribution of the convection term for $f_{r}$ was higher than the $\tilde{P}_{l}$ term. As $T_{i}$ is increased, the location where the kinetic energy of the $f_{r}$ mode is maximum moves further upstream. Therefore a larger proportion of the $\tilde{P}_{l}$ is not captured within the PIV field of view and eventually this missing energy is ‘picked up’ by the convection term. A field of view large enough to capture the full spatial growth and decay of the mode would yield zero total convection. Note that the tip vortices ($3f_{r}$) are formed due to the difference in pressure between the pressure and suction surface of the blades [Andersen et al., 2013]. $3f_{r}$ is therefore energised upstream of the FOV and simply decays (due to different energy transfers) within the FOV under consideration. Interestingly, the $2f_{r}$ mode was found to be energized by the triadic interaction term for all the FST cases in which it could be captured. A similar observation was reported by Biswas and Buxton [2024b] when FST was negligible. $2f_{r}$ was therefore identified as a secondary mode that is produced purely due to non-linear interaction at $\lambda=6$. The magnitude of the triadic interaction term diminished drastically with $T_{i}$ indicating weaker non-linear interaction. For the highest $T_{i}$ cases, the harmonics of $f_{r}$ and $3f_{r}$ were not observed within the field of view and hence no triadic interaction could be captured. The new frequencies $3f_{r}\pm f_{wm}$ for case 2a, although weak, showed a positive contribution from the triadic interaction term, definitively establishing the fact that these are secondary modes that were produced due to the interaction between $3f_{r}$ and $f_{wm}$, similar to those observed in simplified multiscale flows [Baj and Buxton, 2017, Biswas et al., 2022].
@@
Figure 19: (a) Streamwise variation of spanwise averaged mean flow production term for $f_{r}$ for the different FST cases. The variation of $x_{w}$ is shown against (b) $T_{i}$ and (c) $L_{v}$.
@@ 6.1 Triadic interactions
In equation 2, if we only sum over $f_{t}$, we can obtain the triadic energy transfer between $f_{l}$ and $f_{s}$. Biswas and Buxton [2024b] reported these inter-frequency triadic energy transfers in the tip vortex system (formed by the modes $f_{r}-6f_{r}$) for negligible $T_{i}$ and found the same pattern of triadic energy transfers for a range of $\lambda$s ($\lambda\geq 5.5$). Furthermore, the triadic energy transfers were mainly concentrated within the first four frequencies $f_{r}-4f_{r}$. We perform a similar analysis here for $\lambda=6$ for different levels of $T_{i}$. The inter-frequency triadic energy transfers (integrated over the entire domain) are obtained for select cases and shown in figure 18. Note that for the no grid case, the energy exchange pattern is the same as that observed for $\lambda=6$ in [Biswas and Buxton, 2024b]. Interestingly, the same pattern is observed for the low turbulence intensity cases (case 2a). However, the transfers were an order of magnitude weaker as can be seen in figure 18(b) for case 2a. It can also be seen that the new frequencies $3f_{r}\pm f_{wm}$ are energised by $3f_{r}$, i.e. the mode having the higher frequency among the two main interacting modes. This energy pathway is similar to that observed for the secondary modes in the simple multiscale flow discussed in [Biswas et al., 2022]. A similar energy exchange pattern is observed for case 3 (figure 18(c)), but the transfer magnitudes were nearly two orders of magnitude weaker in comparison to the no grid case. These results quantitatively indicate how freestream turbulence weakens the tip vortices which affects the onset of wake recovery.
@@ 6.2 Wake recovery
In figure 19(a), we show the streamwise evolution of $\tilde{P}_{f_{r}}$, integrated along $y$, for four cases of increasing $T_{i}$. Note that, a sign change in $\tilde{P}_{f_{r}}$ is observed even with freestream turbulence. The streamwise location where it occurs moves upstream as freestream turbulence intensity is increased. This clearly shows how the wake recovers more quickly in the presence of freestream turbulence. Figures 19(b-c) show how the wake recovery onset location ($x_{w}$) changed with $T_{i}$ and $L_{v}$ respectively for all the cases. Recently, Biswas and Buxton [2024a] introduced a new length scale (termed as the convective pitch, $Lc$) which was shown to be a better length scale in the near wake compared to $D$. This length scale was defined as $Lc=\pi D/\lambda$ and could be physically interpreted as the distance traveled by a fluid parcel in the freestream in the time taken by the turbine to complete a single revolution. Here $x_{w}$ is presented in terms of $Lc{{}_{eff}}$ which is obtained based on $\lambda_{eff}$ for each case. Note that, for the cases with the highest $T_{i}$s, the $\tilde{P}_{f_{r}}$ term was negligible, at least within the field of view considered, hence, $x_{w}$ could not be measured. Nevertheless, figures 19(b-c) show some interesting trends. Apart from the fact that $x_{w}$ generally reduces with increasing FST, it appears that for low $T_{i}$, $x_{w}$ scales better with $L_{v}$. Although figure 19 lacks data for higher $L_{v}$, it is seen that for the highest $L_{v}$ for which $x_{w}$ could be calculated, $x_{w}$ clearly did not follow the decreasing trend observed for lower $L_{v}$. This indicates that there is likely a threshold length scale for which the initiation of wake recovery is the closest to the turbine. Furthermore, this threshold length scale was close to the average size of the tip vortices near the turbine. The size of the tip vortices was calculated by fitting a circle to the phase averaged vorticity field of the tip vortices ($3f_{r}$) for the no grid case (not shown here). The best fit circle had a diameter $\approx 0.07D$, close to where the trend in figure 19(c) changes. This observation is similar to that of Gambuzza and Ganapathisubramani [2023] who noted that for very large integral length scales in a non-Kolmogorov type flow, the near wake length increased back to a value close to the non-turbulent case. Note that the near wake length is loosely connected to the wake recovery location defined here as both depend on the stability of the tip vortices. Another recent study by Bourhis et al. [2025] looked at the frequency dynamics in the near wake of a model wind turbine exposed to FST. The smallest integral length scales the authors tested were comparable to the largest studied here. Unsurprisingly, they did not find any clear influence of the integral length scales on the strength of the leading tip vortex frequencies.
Figure 20: Regular grids (a) SRG38 and (b) SRG111 and fractal grids (c) SFG5 and (d) SFG7 used in this study (image reproduced from Kankanwadi [2022]).
Figure 21: (a) Mean velocity field and (b) turbulence intensity ($T_{i}$) field behind the regular grid SRG38.
@@ 7 Conclusion
The effect of freestream turbulence (FST) on the multiscale coherent dynamics in a wind turbine wake was studied using particle image velocimetry measurements. A total of 9 FST cases were studied that had varying turbulence intensity ($T_{i}$) and integral length scales ($L_{v}$). The tip vortices were found to break down closer to the turbine in the presence of FST, resulting in an earlier initiation of wake recovery [Lignarolo et al., 2015, Biswas and Buxton, 2024b]. The onset location of wake recovery ($x_{w}$) was found to depend on both $T_{i}$ and $L_{v}$. On average, it reduced monotonically with $T_{i}$. Interestingly, a stronger dependence of $x_{w}$ on $L_{v}$ was observed when $L_{v}$ was smaller than the average size of the tip vortices, likely hinting the presence of a critical $L_{v}$ beyond which the wake recovery onset location can not be pushed further upstream [Gambuzza and Ganapathisubramani, 2023, Bourhis et al., 2025].
The nature of wake meandering with and without FST was elucidated. The frequency spectra in the vicinity of the wake meandering frequency band ($0.15\leq St_{D}\leq 0.4$) changed from a more discrete nature to a more broadband nature as $T_{i}$ was increased. All the FST cases showed stronger wake meandering compared with the no FST case. However, amplitudes of wake meandering did not increase monotonically with the level of FST, highlighting the fact that the full detail of the freestream spectrum is important [Mao and Sørensen, 2018, Gupta and Wan, 2019].
@@
The coherent modes associated with the different frequencies were studied by using Optimal Mode Decomposition (OMD). The OMD algorithm identified a larger number of modes in the vicinity of the wake meandering frequency band that were spatially (and spectrally) similar. These modes were grouped/ clustered by using the mode clustering algorithm recently introduced by Beit-Sadi et al. [2021]. The algorithm significantly simplified the OMD spectra, particularly in the vicinity of the wake meandering frequency band. As the FST level was increased, the tip vortex related modes became weaker and decayed much closer to the turbine. On the other hand, the wake meandering modes for the cases with FST were always found to be more energetic than that for the no grid case.
@@
The energy exchanges to/from the modes were studied using the coherent kinetic energy budget equation derived by Baj and Buxton [2017]. Even with FST, the wake meandering modes were found to be primarily energised by the mean flow, although there were non-negligible contributions from the diffusion term and triadic interactions. In the presence of FST, the wake meandering modes were shown to better utilize the mean velocity gradient, yielding a higher energy transfer from the mean flow. With FST, the pattern of triadic energy exchanges in the tip vortex system remained almost identical to that reported without FST [Biswas and Buxton, 2024b]. However, the magnitudes of the energy exchanges were drastically reduced for elevated FST levels. New secondary frequencies were reported for the first time that were produced through non-linear triadic interaction between the tip vortices and wake meandering induced by FST, similarly to our previous studies on multiscale wakes [Baj and Buxton, 2017, Biswas et al., 2022].
Figure 22: Variation of mean streamwise velocity along the centreline ($y=0$, $z=0$) behind the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The shaded region shows the variation of mean streamwise velocity within the field of view considered.
Figure 23: Variation of turbulence intensity ($T_{i}$) at hub height with distance from the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The dashed vertical lines show the position of the rotor plane in different experiments. The shaded region shows the variation of $T_{i}$ within the field of view considered.
Apendix 1: Grid characterisation
A separate experimental campaign was conducted to study the effect of freestream turbulence on the wind turbine wake. To generate freestream turbulence, 4 different passive turbulence generating grids were used. These are shown in figure 20. Among these two were regular grids and two were fractal grids. The flow fields behind these grids were first characterised using PIV experiments. 2 Phantom v641 cameras were used simultaneously to give a stitched field of view length $\approx 2.45D$ ($49$cm). The FOV was $\approx 18$cm $(0.9D)$ wide and the centre of the FOV coincided with the centre of the flume (i.e. $30$cm from the bottom wall). The grid was moved upstream keeping the optical setup intact and a number of non-synchronous experiments were conducted for a particular grid to obtain time averaged statistics over a certain streamwise distance as required. As an example the mean velocity field and the turbulence intensity field behind the grid SRG38 is shown in figure 21. Note that just behind the grid, there is a significant amount of acceleration. The pressure drop across the grid reduces the water depth slightly, hence continuity yields the acceleration. Although the acceleration reduces with downstream distance, the local mean velocity is slightly greater than $U_{\infty}$ even at the farthest location measured. This increment in the freestream velocity needs to be taken into account as it can result in slight changes to the tip speed ratio of the turbine. The turbulence intensity is the highest ($\approx 14\%$) near the grid and monotonically reduces to $\approx 3.5\%$ further downstream.
The variation of the mean streamwise velocity behind all the grids is shown in figure 22. The solid black line shows the variation of the mean centreline velocity (along $y=0,z=0$). The jumps in the variation (especially for SFG7) are due to the fact that mean data from multiple non-synchronous PIV experiments were stitched together. The streamwise mean velocity also varies along the $y$ direction. Therefore, the transverse variation of the mean velocity is indicated by the shaded region. The grid is located at $x/D=0$. The dashed vertical lines show the position of the wind turbine in different experiments. Similarly, the variation of $T_{i}$ behind the grids is shown in figure 23.
Figure 24: Streamwise autocorrelation function of transverse velocity fluctuation (a) $2D$ and (b) $5D$ downstream of the grid SRG38 at the geometric centreline ($y=0,z=0$). In (b), the black curve shows the fitted bi-Gaussian curve. The strongest secondary autocorrelation peak is shown by the red ‘+’ sign. The red dashed line shows the cutoff, the data below which was discarded during fitting.
Figure 25: Variation of integral length scale ($L_{v}$) at hub height with distance from the grids (a) SRG38, (b) SRG111, (c) SFG7 and (d) SFG5. The dashed vertical lines show the position of the rotor plane in different experiments. The shaded region shows the variation of $L_{v}$ within the field of view considered.
Figure 26: The nature of the different auto-corrletaion functions (a) $2.5D$ downstream of the grid SRG38 and (b) $14D$ downstream of the grid SFG5 at the geometrical centreline ($y=0,z=0$).
Another important quantity to characterise the nature of freestream turbulence is the integral length scale ($L_{v}$). $L_{v}$ is estimated using equation 3.

### v2

@@ 5 Energy exchanges
Biswas and Buxton [2024b] showed that in the presence of negligible freestream turbulence, the wake meandering mode acted like a ‘primary mode’, extracting energy from the mean flow. In the presence of FST, the mean flow production ($\tilde{P}_{l}$) term is still found to be the dominant energy source for wake meandering. The $\tilde{P}_{l}$ terms for the dominant wake meandering modes in the presence of FST were always found to be stronger than that for the no grid case which is in line with the discussion in section 3. Furthermore, the dominant wake meandering modes for case 2c drew the highest amount of energy from the mean flow. This is in line with the highest far wake meandering amplitude seen for this FST case in figure 13. The $\tilde{P}_{l}$ term then became relatively weaker as $T_{i}$ was increased further, likely due to the wake meandering modes becoming less coherent (as observed in figure 17) and less efficient at extracting energy from the mean flow. This was accompanied by an enhancement in the relative contributions from the triadic interaction and diffusion term for the highest $T_{i}$ case.
@@
The tip vortex related modes in the presence of FST showed energy exchanges similar to the no grid case, however, the magnitudes of the terms decreased as $T_{i}$ was increased. $f_{r}$ for the no grid case was primarily energized by the mean flow production ($\tilde{P}_{l}$) term. For the FST cases, the relative contribution of the convection term for $f_{r}$ was higher than the $\tilde{P}_{l}$ term. As $T_{i}$ is increased, the location where the kinetic energy of the $f_{r}$ mode is maximum moves further upstream. Therefore a larger proportion of the $\tilde{P}_{l}$ is not captured within the PIV field of view and eventually this missing energy is ‘picked up’ by the convection term. A field of view large enough to capture the full spatial growth and decay of the mode would yield zero total convection. Note that the tip vortices ($3f_{r}$) are formed due to the difference in pressure between the pressure and suction surface of the blades [Andersen et al., 2013]. $3f_{r}$ is therefore energised upstream of the FOV and simply decays (due to different energy transfers) within the FOV under consideration. Interestingly, the $2f_{r}$ mode was found to be energized by the triadic interaction term for all the FST cases in which it could be captured. A similar observation was reported by Biswas and Buxton [2024b] when FST was negligible. $2f_{r}$ was therefore identified as a secondary mode that is produced purely due to non-linear interaction at $\lambda=6$. The magnitude of the triadic interaction term diminished drastically with $T_{i}$ indicating weaker non-linear interaction. For the highest $T_{i}$ cases, the harmonics of $f_{r}$ and $3f_{r}$ were not observed within the field of view and hence no triadic interaction could be captured.
The intriguing question is why is the $\tilde{P}_{l}$ term at all stronger for the wake meandering modes in the presence of FST? To answer this we need to look at the composition of the $\tilde{P}_{l}$ term. As discussed in Biswas and Buxton [2024b], utilizing the fact that the transverse gradient of streamwise mean velocity ($\frac{\partial\bar{u}}{\partial y}$) is much stronger than the other mean velocity gradients and assuming that the velocity components of the different modes are uncorrelated ($\overline{\tilde{u}^{l}\tilde{v}^{n}}\approx 0$), we can approximate the mean flow production term as $\tilde{P}_{l}\approx-\overline{\tilde{u}^{l}\tilde{v}^{l}}\frac{\partial\overline{u}}{\partial y}$. This was verified to be true for all the modes and is not shown here for brevity, a more detailed discussion can be found in Biswas and Buxton [2024b]. In figure 21, the profiles of $\overline{\tilde{u}^{l}\tilde{v}^{l}}$ and $\frac{\partial\overline{u}}{\partial y}$ are shown for the two important wake meandering modes (as shown in figure 17) for the four cases (1, 2c, 3b and 4b) at two streamwise locations, $x/D=1$ (representative of near wake) and $x/D=4$ (representative of far wake). Note that, the no grid case shows the strongest gradients in the tip shear layer. However the wake meandering modes remain sheltered well within the shear layer (see figure 17(a,b)) in the near wake as well as the far wake, inhibiting it from utilizing the stronger gradients available in the shear layer. For the other cases with FST, in the near wake, the wake meandering modes still remain sheltered, explaining the similar wake meandering amplitudes observed for all the cases in the near wake (figure 13). Further downstream, due to the earlier breakdown of the tip vortices, the modes extend beyond the shear layer (see figure 17). As a result, although the mean velocity gradients are slightly weaker (due to enhanced turbulent mixing), the modes do a better job of exploiting this shear thereby yielding a higher rate of energy transfer from the mean flow.
Figure 22: Inter frequency triadic energy transfers for (a) no grid case, (b) case 2a, (c) case 2c and (d) case 3b.
@@
Figure 23: (a) Streamwise variation of spanwise averaged mean flow production term for $f_{r}$ for the different FST cases. The variation of the wake recovery onset location ($x_{w}$) is shown against (b) $T_{i}$ and (c) $L_{v}$. $x_{w}$ is scaled by the convective length scale ($Lc_{eff}=\pi D/\lambda_{eff}$).
@@ 5.1 Triadic interactions
[…]
@@ 5.2 Wake recovery
In figure 23(a), we show the streamwise evolution of $\tilde{P}_{f_{r}}$, summed along $y$, for four cases of increasing $T_{i}$. Note that, a sign change in $\tilde{P}_{f_{r}}$ is observed even with freestream turbulence. The streamwise location where it occurs one an average moves upstream as freestream turbulence intensity is increased. This clearly shows how the wake recovers more quickly in the presence of freestream turbulence. Figures 23(b-c) show how the wake recovery onset location ($x_{w}$) changed with $T_{i}$ and $L_{v}$ respectively for all the cases for which it could be measured. For the cases with the highest $T_{i}$s, the $\tilde{P}_{f_{r}}$ term was negligible, at least within the field of view considered, hence, $x_{w}$ could not be measured. Recently, Biswas and Buxton [2024a] introduced a new length scale (termed as the convective pitch, $Lc=\pi D/\lambda$) which could be physically interpreted as the distance traveled by a fluid parcel in the freestream in the time taken by the turbine to complete a single revolution. Note that $Lc$ is not exactly the same as the pitch of the tip vortices as the tip vortices propagate at a velocity slightly smaller than the freestream velocity. In figure 23(b-c), $x_{w}$ is presented in terms of $Lc{{}_{eff}}$ which is obtained based on $\lambda_{eff}$ for each case. Apart from the fact that $x_{w}$ generally reduces with increasing FST, it appears that for low $T_{i}$, $x_{w}$ scales better with $L_{v}$. Although figure 23 lacks data for higher $L_{v}$, it is seen that for the highest $L_{v}$ for which $x_{w}$ could be calculated, $x_{w}$ clearly did not follow the decreasing trend observed for lower $L_{v}$. This indicates that there is could be a threshold length scale for which the initiation of wake recovery is the closest to the turbine. Furthermore, this threshold length scale was close to the average size of the tip vortices near the turbine. The size of the tip vortices was calculated by fitting a circle to the phase averaged vorticity field of the tip vortices ($3f_{r}$) for the no grid case (not shown here). The best fit circle had a diameter $\approx 0.07D$, close to where the trend in figure 23(c) changes. This observation is similar to that of Gambuzza and Ganapathisubramani [2023] who noted that for very large integral length scales in a non-Kolmogorov type flow, the near wake length increased back to a value close to the non-turbulent case. Note that the near wake length is loosely connected to the wake recovery location defined here as both depend on the stability of the tip vortices. Another recent study by Bourhis et al. [2025] looked at the frequency dynamics in the near wake of a model wind turbine exposed to FST. The smallest integral length scales the authors tested were comparable to the largest studied here. Unsurprisingly, they did not find any clear influence of the integral length scales on the strength of the leading tip vortex frequencies.
@@ 6 Conclusion
The effect of freestream turbulence (FST) on the multiscale coherent dynamics in a wind turbine wake was studied using particle image velocimetry measurements. A total of 10 FST cases were studied that had varying turbulence intensity ($T_{i}$) and integral length scales ($L_{v}$) much smaller than the turbine diameter. The tip vortices were found to break down closer to the turbine in the presence of FST, resulting in an earlier initiation of wake recovery. The onset location of wake recovery ($x_{w}$) was found to move closer to the turbine as $T_{i}$ and $L_{v}$ were increased. A stronger dependence of $x_{w}$ on $L_{v}$ was observed, at least for the low $T_{i}$ cases considered.
The nature and amplitude of wake meandering at varying FST levels was elucidated. For all the cases, low frequency peaks were observed in the wake that fell in the generally accepted wake meandering frequency band ($0.15\leq St_{D}\leq 0.4$). In the near wake, the power associated with wake meandering was more comparable for all the cases, showing no clear trend with either $T_{i}$ or $L_{v}$. This indicated that the process of initiation of wake recovery was the same for all the cases, likely related to a turbine instability [Howard et al., 2015, Foti et al., 2016, Biswas and Buxton, 2024a]. In the far wake however, as the wake meandering power for the no grid case (negligible FST) decayed, that for the cases with non-negligible FST levels sustained, primarily due to the earlier breakdown of the tip vortices. For the highest FST cases a broadening of the wake meandering frequency band was observed [Bourhis et al., 2025].
@@
The coherent modes associated with the different frequencies were studied by using Optimal Mode Decomposition (OMD). The OMD algorithm identified a larger number of modes in the vicinity of the wake meandering frequency band that were spatially (and spectrally) similar. These modes were grouped/ clustered by using a slightly modified version of mode clustering algorithm originally introduced by Beit-Sadi et al. [2021]. The algorithm significantly simplified the OMD spectra, particularly in the vicinity of the wake meandering frequency band. As the FST level was increased, the tip vortex related modes became weaker and decayed much closer to the turbine. On the other hand, the wake meandering modes for the cases with FST showed a larger spanwise width.
@@
The energy exchanges to/from the modes were studied using the coherent kinetic energy budget equation derived by Baj and Buxton [2017]. Even with FST, the wake meandering modes were found to be primarily energised by the mean flow, although there were non-negligible contributions from the diffusion term and triadic interactions. In the presence of FST, the earlier breakdown of the tip vortices allowed the wake meandering modes to better utilize the mean velocity gradient, yielding a higher energy transfer from the mean flow. With FST, the pattern of triadic energy exchanges in the tip vortex system remained almost identical to that reported with negligible FST [Biswas and Buxton, 2024b]. However, the magnitudes of the energy exchanges were drastically reduced for elevated FST levels. New secondary frequencies were reported for the first time that were produced through non-linear triadic interaction between the tip vortices and wake meandering induced by FST, similarly to our previous studies on multiscale wakes [Baj and Buxton, 2017, Biswas et al., 2022].
Our study confirms that large-scale turbulence ($L_{v}>D$) is not compulsory to have energetic modes in the meandering frequency band as assumed in the dynamic wake meandering model (DWM). When wake meandering is primarily driven by the turbine, even small scale turbulence can significantly alter wake meandering dynamics, indirectly through breaking down the shield of the tip vortices. The breakdown of the tip vortices introduced several new frequencies in the tip shear layer making it more broadband and turbulent like. The interaction between the shear layer and the wake meandering modes after the tip vortices have broken down is not understood properly. In addition, this work considered the wake only up to a distance of $x=5D$. Future work should conduct a similar analysis looking at the energy exchanges associated wake meandering further downstream. A pertinent question to answer would be: can wake meandering turn into a completely stochastic process further downstream although its initiation could be related to an intrinsic instability of the turbine? In the far wake, possible links could be found between wake meandering and meandering observed in other flows [Bailey et al., 2018, Bölle, 2024].
Apendix 1: Uncertainties in the estimation of the integral length scales
Several factors contribute to the uncertainties in the estimation of the integral length scales. Firstly the correlation functions themselves may depend on the length of the time series used. To check this, we evaluated the $R_{12}^{{}^{\prime}}$ for data truncated in time. Figure 24 here shows the variation of $R_{12}^{{}^{\prime}}$ at the hub height for two different cases, 2c and 4b with different data length (here $N=5456$ is the total data length in time). $R_{12}^{{}^{\prime}}$ is found to be generally well converged when the full data is used. The uncertainties in the estimation of the integral length scales were introduced mainly due to the inherent inhomogeneity of the flow cases and due some extent to the use of a fitting function (for cases where it did not fall off to zero).
Figure 24: Effect of the data length on the estimation of $R_{12}^{{}^{\prime}}$ for the cases 2c and 4b.
Appendix 2: OMD based mode clustering
Dynamic Mode Decomposition (DMD) assumes that there is a linear dependency between two successive snapshots ($\boldsymbol{x_{n}},\boldsymbol{x_{n+1}}$) of the flow which remains constant over time. This can be written as:
$\boldsymbol{x_{n+1}}=\boldsymbol{Ax_{n}},\hskip 28.45274pt\boldsymbol{A}\in\mathbb{R}^{p\times p}$ (A2.4)
Here $p$ denotes the length of a snapshot (the number of PIV vectors). DMD finds $\boldsymbol{A}$ by solving the minimization problem:
$\min\Big\|[\boldsymbol{v_{2}},...\boldsymbol{v_{N}}]-\boldsymbol{A}[\boldsymbol{v_{1}},...\boldsymbol{v_{N-1}}]\Big\|^{2}$ (A2.5)
Here, $N$ is the number of PIV snapshots. In Optimal Mode Decomposition (OMD), $\boldsymbol{A}$ is represented in a rank constrained form ($\boldsymbol{LML^{T}}$) and a two-way optimization is performed (with optimization variables $\boldsymbol{L}$ and $\boldsymbol{M}$):
$\min\Big\|[\boldsymbol{v_{2}},...\boldsymbol{v_{N}}]-\boldsymbol{LML^{T}}[\boldsymbol{v_{1}},...\boldsymbol{v_{N-1}}]\Big\|^{2}$ (A2.6)
Here, $\boldsymbol{L}$ represents the low order subspace of the flow field and $\boldsymbol{M}$ represents the dynamics of $\boldsymbol{L}$. Wynn et al. [2013] showed that in DMD, $\boldsymbol{L}$ was taken as the POD subspace of the flow. Thereafter, the best dynamical representation ($\boldsymbol{M}$) of the flow on that subspace was obtained. Wynn et al. [2013] argued that the POD subspace is not necessarily the best choice for $\boldsymbol{L}$, hence suggesting a more general, two-way optimization. Finally, after finding the eigenvectors ($\boldsymbol{z_{m}}$) and eigenvalues ($\lambda_{M}$}) of $\boldsymbol{M}$, the OMD eigenvectors ($\boldsymbol{\phi_{OMD}}$) and eigenvalues ($\lambda_{OMD}$}) are obtained as:
$\boldsymbol{\phi_{OMD}}=\boldsymbol{Lz_{M}},\hskip 28.45274pt{\lambda_{OMD}}=\log({\lambda_{M}})/{\Delta}t$ (A2.7)